Solve
An experienced accountant can balance the books twic as fast as a new accountant, Working together it take the accountants 10 hours. How long would it take the experienced accountant workign alone?

Answer by checkley77(12844)

THE FORMULA FOR THESE TYPE OF WORK TOGETHER PROBLEMS IS:
XY/(X+Y)=TOTAL TIME
HERE WE HAVE X & 2X FOR THE TWO TIMES & 10 HOURS FOR THE TOTAL TIME.
X*2X/(X+2X)=10
2X^2/3X=10
2X^2=10*3X
2X^2=30X
2X^2-30X=0
2X(X-15)=0
X-15=0
X=15 HOURS FOR THE EXPERIENCED ACCOUNTANT.
2*15=30 HOURS FOR THE NEW ACCOUNTANT.
PROOF:
15*30/(15+30)=10
450/45=10
10=10